# Beam-Beam Backgrounds for the Cool Copper Collider

### v1

@@ 5.3 Spatial distribution, detector impact, and mitigation strategies
where the number of hits in each subdetector $D$ can be calculated by multiplying the values of Table˜7 in Appendix˜C with the bunch train duration for each scenario, and the number of cells in $D$ is obtained from the detector geometry outlined in Section˜3.1 and assuming the cell sizes of Table˜3. Finally, $S_{D}$ is a safety factor used to account for potential occupancy increase due to mismodeling effects, both at the background generation and the detector simulation level, and $C_{D}$ is a cluster size factor to account for charge sharing effects. For the purposes of this study, and following similar assumptions previously made in the literature [40, 41], we adopt $S_{D}=2$ for all subdetectors, and $C_{D}=3$ for the vertex detector, with $C_{D}=1$ everywhere else [^4] For MAPS with $\simeq 10$ µm pitch operated at nominal thresholds, we assume a mean cluster size of 3, consistent with ALPIDE beam-test characterizations [42, 43]. For all other subdetectors, no charge sharing is assumed..
The average occupancy for the vertex and tracker detectors are gathered in Table˜4 for the various C3 parameter scenarios. Following previous ILC studies [44], an occupancy threshold of $10^{-4}$ is defined as the design target for the tracking detectors. This threshold preserves hit efficiency at levels not significantly lower than the intrinsic sensor efficiency, thereby maintaining robust tracking performance essential for precision vertexing and heavy-flavor tagging. We notice that the average occupancy for the vertex exceeds $10^{-4}$ for all C3 operating scenarios, rising to $\mathscr{O}(10^{-3})$ for the highest-luminosity scenario at 550 GeV. On the contrary, the average occupancy for the tracker barrel and endcap detectors remains well below the $10^{-4}$ limit. With respect to the baseline C3 scenario at 250 GeV, the occupancy increases approximately by two (five)-fold for the s.u. (high-$\mathscr{L}$) configuration at 250 GeV, whereas for the higher energy runs at 550 GeV, we observe on average a two-, four-, and eight-fold increase for the baseline, s.u. and high-$\mathscr{L}$ scenarios, respectively.
@@
These occupancy results reveal that the tracker barrel and endcap remain within occupancy limits without multi-buffer operation. The vertex detector, while most exposed to beam–beam backgrounds, requires only modest local buffer depths, two to three. This could be addressed by employing a dedicated readout with on-pixel memory capable of storing two to three hits locally during the bunch train. This approach would provide sufficient buffering to maintain occupancies below $10^{-4}$ throughout the entire tracking system and across all C3 scenarios, thus ensuring that all detector hits can be successfully recorded. This is within limits of the buffer depth of four previously found to be necessary to ensure that the maximum layer occupancy for the SiD vertex barrel does not exceed the occupancy threshold under the ILC beam conditions [44].
@@ Appendix A Additional information for the SiD detector concept
Table˜6 reproduces the key radii, longitudinal extents, and technologies of the various SiD subsystems. Relative to the Detailed Baseline Design [32], the main functional difference is the transition from Resistive Plate Chambers (RPCs) to scintillator tiles read out by Silicon photomultiplier (SiPMs) in the HCAL; all absorber thicknesses and support material fractions follow the updated study of  [45].
Table 6: Main parameters of the SiD concept. The changes with respect to the baseline SiD design [32] are explained in the text.
@@
For the vertex system specifically, SiD employs a barrel-disk geometry with three functionally distinct regions, illustrated in Fig.˜12. The five-layer vertex barrel provides precise tracking in the central region, while four vertex endcap disks at intermediate $z$ ensure hermetic coverage and smooth pattern recognition in the transition to forward angles. Three additional forward vertex disks extend coverage to $\cos\theta\approx 0.984$ [32], bridging the gap to the tracker.

### v2

@@ 5.3 Spatial distribution, detector impact, and mitigation strategies
where the number of hits in each subdetector $D$ can be calculated by multiplying the values of Table˜7 in Appendix˜C with the bunch train duration for each scenario, and the number of cells in $D$ is obtained from the detector geometry outlined in Section˜3.1 and assuming the cell sizes of Table˜3. Finally, $S_{D}$ is a safety factor used to account for potential occupancy increase due to mismodeling effects, both at the background generation and the detector simulation level, and $C_{D}$ is a cluster size factor to account for charge sharing effects. For the purposes of this study, and following similar assumptions previously made in the literature [Arominski:2704642, Boscolo:2023grv], we adopt $S_{D}=2$ for all subdetectors, and $C_{D}=3$ for the vertex detector, with $C_{D}=1$ everywhere else [^4] For MAPS with $\simeq 10$ µm pitch operated at nominal thresholds, we assume a mean cluster size of 3, consistent with ALPIDE beam-test characterizations [Rinella:2022htp, Andronic:2025hcm]. For all other subdetectors, no charge sharing is assumed..
The average occupancy for the vertex and tracker detectors are gathered in Table˜4 for the various C3 parameter scenarios. Following previous ILC studies [Schutz:2018ynd, behnke2013internationallinearcollidertechnical], an occupancy threshold of $10^{-4}$ is defined as the design target for the tracking detectors. This threshold has been validated through extensive simulation studies for the ILC detector concepts and was found to preserve hit efficiency at levels not significantly lower than the intrinsic sensor efficiency, thereby maintaining robust tracking performance essential for precision vertexing and heavy-flavor tagging. It thus serves as a practical benchmark to enable direct comparison with other linear collider background studies.
Using this threshold, we notice that the average occupancy for the vertex exceeds $10^{-4}$ for all C3 operating scenarios, rising to $\mathscr{O}(10^{-3})$ for the highest-luminosity scenario at 550 GeV. On the contrary, the average occupancy for the tracker barrel and endcap detectors remains well below the $10^{-4}$ limit. With respect to the baseline C3 scenario at 250 GeV, the occupancy increases approximately by two (five)-fold for the s.u. (high-$\mathscr{L}$) configuration at 250 GeV, whereas for the higher energy runs at 550 GeV, we observe on average a two-, four-, and eight-fold increase for the baseline, s.u. and high-$\mathscr{L}$ scenarios, respectively.
@@
These occupancy results reveal that the tracker barrel and endcap remain within occupancy limits without multi-buffer operation. The vertex detector, while most exposed to beam–beam backgrounds, requires only modest local buffer depths, two to three. This could be addressed by employing a dedicated readout with on-pixel memory capable of storing two to three hits locally during the bunch train. This approach would provide sufficient buffering to maintain occupancies below $10^{-4}$ throughout the entire tracking system and across all C3 scenarios, thus ensuring that all detector hits can be successfully recorded. This is within limits of the buffer depth of four previously found to be necessary to ensure that the maximum layer occupancy for the SiD vertex barrel does not exceed the occupancy threshold under the ILC beam conditions [Schutz:2018ynd].
@@ Appendix A Additional information for the SiD detector concept
Table˜6 reproduces the key radii, longitudinal extents, and technologies of the various SiD subsystems. Relative to the Detailed Baseline Design [behnke2013internationallinearcollidertechnical], the main functional difference is the transition from Resistive Plate Chambers (RPCs) to scintillator tiles read out by Silicon photomultiplier (SiPMs) in the HCAL; all absorber thicknesses and support material fractions follow the updated study of  [Breidenbach:2021sdo].
Table 6: Main parameters of the SiD concept. The changes with respect to the baseline SiD design [behnke2013internationallinearcollidertechnical] are explained in the text.
@@
For the vertex system specifically, SiD employs a barrel-disk geometry with three functionally distinct regions, illustrated in Fig.˜12. The five-layer vertex barrel provides precise tracking in the central region, while four vertex endcap disks at intermediate $z$ ensure hermetic coverage and smooth pattern recognition in the transition to forward angles. Three additional forward vertex disks extend coverage to $\cos\theta\approx 0.984$ [behnke2013internationallinearcollidertechnical], bridging the gap to the tracker.